OpenStudy (anonymous):

an arc of length 12 meters is formed by a central angle A on a circle of r=9 the measure of A in degrees(2 decimal places() is a. 42.97 b. 8.38 c. 76.39 d. 1.30

OpenStudy (anonymous):

b maybe?

OpenStudy (anonymous):

\[s = r \theta\] Solve for theta. \[\theta = s/r\] \[\theta = 1.3333\] So your answer is D

OpenStudy (anonymous):

it was c

OpenStudy (anonymous):

First, find the circumference of the circle. Then, find the ratio of arc to circumference. Then, use this ratio against 360 degrees to find the angle A.

OpenStudy (anonymous):

circumference of circle = 2*p1*9 = 18 pi required angle = 12* 360 ------- 18 pi = 240 / pi = 76.39 degrees